OpenStudy (anonymous):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now? 15 10 7 5

OpenStudy (anonymous):

if sam is now x then his coach x+3 before 5 year sam x-5 and its coach=5(x-5) so 5(x-5)=x+3-5 4x=23 x=23/4

OpenStudy (anonymous):

Solve the following for coach and Sam:\[{coach = 3 Sam, coach - 5 = 5 (Sam - 5)}\]

OpenStudy (anonymous):

\[\{\text{coach}\to 30,\text{Sam}\to 10\} \]

Parth (parthkohli):

s - Sam 3s - Coach\[\rm 5(s - 5) = 3s - 5\]

Parth (parthkohli):

@robtobey Top tip: use \rm to turn the text into text and still be able to use other \(\rm \LaTeX\) commands. See this one:\[\rm \{coach\rightarrow 30,Sam\rightarrow 10\}\]

OpenStudy (anonymous):

@ParthKohli Used Mathematica to do the calculation. This program returns answers as replacement rules. Normally I copy and paste out the right arrows with "=" signs by hand. Sorry about that.